OpenStudy (anonymous):

can someone help... 1. (7)/(1-√3) A. -(7-7√2)/(2 B. -(7+7√2)/(2) C. (4+6√3)/(3) D (2+7√3)/(2)

OpenStudy (anonymous):

@Beautiful_lier96 do u know how to rationalize?

OpenStudy (anonymous):

kinda..

OpenStudy (anonymous):

@niksva i know the bottom number is -2

OpenStudy (anonymous):

u r right what do u get in numerator?

OpenStudy (anonymous):

umm 7+√3..... so im guessing it would be letter B

OpenStudy (anonymous):

check again u will get { 7+ 7 sqrt(3) }

OpenStudy (anonymous):

how ? n the answer isnt b ?

OpenStudy (anonymous):

i think there is something wrong in the options it cant be b

geerky42 (geerky42):

\[\dfrac{7}{1-\sqrt3} = \dfrac{7}{1-\sqrt3} \cdot \dfrac{1+\sqrt3}{1+\sqrt3} \]

OpenStudy (anonymous):

the answer choice for d is (2+7√3)/2

OpenStudy (anonymous):

it cannot be d

geerky42 (geerky42):

It is B, but instead of √2, it should be √3 Perhaps it's a typo?

OpenStudy (anonymous):

I think, by mistake, someone has wrote sqrt(2) in options a and b in place of sqrt(3)

geerky42 (geerky42):

\(\Huge \text{^}\text{^}\text{^}\text{^}\text{^}\text{^}\text{^}\text{^}\text{^}\text{^}\text{^}\text{^}\text{^}\text{^}\text{^}\text{^}\text{^}\text{^}\text{^}\text{^}\text{^}\text{^}\text{^}\text{^}\text{^}\text{^}\text{^}\text{^}\text{^}\text{^}\)

OpenStudy (anonymous):

according to me , answer is - (7+7√3)/2

geerky42 (geerky42):

Same here.

OpenStudy (anonymous):

i just put it as be. can you help with another question or 2 ? 1. (5√32)-(7√8) A. 4√2 B. 3√2 C. 6√3 D. 6√2 2. (2√15xy³)(3√30x³y²) A. 4x²y³√5xy B. 3x²√2y C. 90x²y²√2y D. x²y²√xy